# VL-ICL Bench: The Devil in the Details of Benchmarking Multimodal In-Context Learning

### v1

@@ Abstract
Large language models (LLMs) famously exhibit emergent in-context learning (ICL) – the ability to rapidly adapt to new tasks using few-shot examples provided as a prompt, without updating the model’s weights. Built on top of LLMs, vision large language models (VLLMs) have advanced significantly in areas such as recognition, reasoning, and grounding. However, investigations into multimodal ICL have predominantly focused on few-shot visual question answering (VQA), and image captioning, which we will show neither exploit the strengths of ICL, nor test its limitations. The broader capabilities and limitations of multimodal ICL remain under-explored. In this study, we introduce a comprehensive benchmark VL-ICL Bench for multimodal in-context learning, encompassing a broad spectrum of tasks that involve both images and text as inputs and outputs, and different types of challenges, from perception to reasoning and long context length. We evaluate the abilities of state-of-the-art VLLMs against this benchmark suite, revealing their diverse strengths and weaknesses, and showing that even the most advanced models, such as GPT-4, find the tasks challenging. By highlighting a range of new ICL tasks, and the associated strengths and limitations of existing models, we hope that our dataset will inspire future work on enhancing the in-context learning capabilities of VLLMs, as well as inspire new applications that leverage VLLM ICL. The code and dataset are available at https://github.com/ys-zong/VL-ICL.
@@ 1 Introduction
With the scaling of model size, large language models (LLMs) famously exhibit the emergent capability of in-context learning (ICL) (Brown et al., 2020; Dong et al., 2022). This refers to the ability to learn from analogy within a single feed-forward pass – thus, enabling the model to learn completely new tasks using a few input-output examples, without requiring any updates to the model parameters. This training-free nature of ICL has led to its rapid and broad application across a wide range of scenarios and applications, as illustrated by benchmarks such as (Hendrycks et al., 2021; Zhong et al., 2023; Srivastava et al., 2023; Cobbe et al., 2021).
@@
Vision large language models (VLLMs) are typically built on a base LLM, by augmenting it with a vision encoder and/or decoder connected by some stitching mechanism (Liu et al., 2023a, b; Bai et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Koh et al., 2023; Ge et al., 2024). These models have rapidly advanced alongside LLMs, and attracted significant attention for their remarkable multimodal capabilities in zero-shot recognition, reasoning, grounding, and visual question answering (VQA) among other capabilities. These capabilities have been thoroughly tested by a range of recent benchmark suites (Liu et al., 2023c; Fu et al., 2023; Li et al., 2023c; Yu et al., 2023). Meanwhile, VLLMs are also widely presumed to inherit in-context learning (ICL) capabilities from their base LLM. However, their abilities in this respect are poorly evaluated and poorly understood. Current VLLMs studies mainly report their zero-shot capabilities measured by the benchmarks above, while ICL is usually only evaluated qualitatively, or as a secondary consideration via few-shot visual question answering (VQA) or image captioning (Bai et al., 2023; Awadalla et al., 2023; Sun et al., 2023a; Laurençon et al., 2023), with a notable deficiency in quantitative assessment across a wider spectrum of ICL tasks. This is presumably due to the ready availability of VQA and captioning benchmarking infrastructure. However, we will show that captioning and VQA tasks are not ideal for ICL evaluation: They neither truly exploit the ability of ICL to improve performance from examples; nor they test the limits of what ICL can do, in order to motivate future VLLM research to better exploit and expose the underpinning LLM’s ICL ability.
@@
Figure 1: Illustration of the different tasks in VL-ICL Bench. Image-to-text tasks are in the first three rows, while text-to-image tasks are in the bottom row. Image-to-text tasks in the third row do reasoning on interleaved image-text inputs.
@@
To enhance the understanding of multimodal ICL and assess the ICL capabilities of state-of-the-art VLLMs, we introduce a novel benchmark suite VL-ICL Bench (Figure 1), tailored for assessing VLLM in-context learning. Our benchmark suite incorporates both text-output and image-output tasks, and is designed to test various facets of VLLMs, including fine-grained perception, reasoning, rule induction, and context-length. We conduct comprehensive evaluations of state-of-the-art VLLMs that are capable of processing interleaved image-text as inputs on our benchmark. Results reveal that although certain models exhibit reasonable performance on specific tasks, none demonstrate uniform excellence across the entire spectrum of tasks, and some models perform near chance level on some tasks. We hope that this systematic study of different opportunities and challenges for multimodal ICL will support practitioners to know what is currently possible and impossible in terms of training-free learning of new multimodal tasks, and spur VLLM model developers to study how to expose as much as possible of the LLM’s ICL ability to the multimodal world.
@@
To summarise our contributions: (1) We demonstrate the limitations inherent in the common practice of quantitatively evaluating VLLM ICL via VQA and captioning. (2) We introduce the first thorough and integrated benchmark suite of ICL tasks covering diverse challenges including perception, reasoning, rule-induction, long context-length and text-to-image/image-to-text. (3) We rigorously evaluate a range of state-of-the-art VLLMs on our benchmark suite, and highlight their diverse strengths and weaknesses, as well the varying maturity of solutions to different ICL challenges.
@@ 2.2 Common Practice in ICL Evaluation
The benchmarks that have been most popular in prior attempts at quantitative evaluation of multimodal ICL are VQA and captioning. We focus our discussion in this section on image-to-text models (Alayrac et al., 2022; Bai et al., 2023; Li et al., 2023b; Laurençon et al., 2023; Awadalla et al., 2023), as in-context text-to-image models (Ge et al., 2024; Koh et al., 2023) are relatively less common and less mature, so there is no common evaluation practice yet. In the case of captioning, the context set $S$ contains examples of images $x$ and captions $y$; while for VQA the context $S$ contains image-question pairs $x$ and answers $y$.
@@
Figure 2(a) plots the ICL performance of six state-of-the-art VLLMs on three popular benchmarks – MathVista VQA (Lu et al., 2024), VizWiz VQA (Gurari et al., 2018), and COCO captioning (Lin et al., 2014) for varying numbers of training examples (shots). While the performance of the different models varies, the key observation is that most of the lines shown are only weakly increasing. It means that for all models, there is limited improvement achieved by ICL (shots $>0$) compared to the zero-shot case (shots $=0$). This is because, while the context set $S$ illustrates the notion of asking and answering a question or captioning images, the baseline VLLM $\theta$ is already quite good at VQA and captioning. The limiting factors in VLLM captioning and VQA are aspects such as detailed perception, common sense knowledge, etc. – all of which are fundamental challenges to the VLLM, and not aspects that can reasonably be taught by a few-shot support set.
@@
Given the discussion above, it is unclear why performance should improve with shots at all? We conjectured that this is largely due to the VLLM learning about each dataset’s preferred answer style, rather than learning to better solve multimodal inference tasks per-se. For example, in captioning zero-shot VLLMs tend to produce more verbose captions than the ground truth in COCO, and they learn to be more concise through ICL. Meanwhile, for VQA, there is a standard practice of evaluating based on string match between the ground-truth answer and the model-provided answer. For example, VizWiz has unanswerable questions, which some VLLMs answer with ‘‘I don’t know’’ which would not be a string matched against a ground truth ‘‘Unanswerable’’. Some models thus learn about answer-formatting (e.g., preferred terminology; avoid using any preface or postface that may not satisfy a string match) from the context set. This is indeed a kind of ICL, but perhaps not what one expects to be learning in VQA. To validate this conjecture, we repeat the previous evaluation, but using soft matching to eliminate the impact of answer format learning. For VQA, we use a pretrained LLM to determine whether the prediction is semantically equivalent to the ground-truth while for captioning, we use an LLM to score the quality of the generated caption on a scale of score 1-10 (details in the Appendix). Fig. 2(b) shows that the curves have almost fully flattened out, with zero-shot performance having improved. Fig. 2(c) quantifies this difference by showing the average rate of improvement with shots for exact match and LLM match. The change to LLM validation almost completely eliminates any benefit of ICL over zero-shot.
@@
In contrast to the above, popular LLM ICL benchmarks in the language domain do usually exhibit non-trivial ICL learning (Brown et al., 2020; Dong et al., 2022). Figure 3 shows three state-of-the-art VLLMs along with their corresponding base LLMs, evaluated on three popular NLP tasks (AGNews (Zhang et al., 2015), MIT Movies (Ushio and Camacho-Collados, 2021) and TREC (Voorhees and Tice, 2000)). We can see that in contrast to the VQA/captioning benchmarks, models’ zero-shot performance is often substantially improved by few-shot ICL. This result confirms that the LLM components in VLLMs do inherit the ICL ability of their base LLM. However, it raises the question of how we can meaningfully exploit and measure the ICL ability of VLLMs in the multimodal context. In the next section, we introduce our benchmark VL-ICL Bench, which does exactly this.
(a) Standard metrics for evaluation.
@@
Figure 2: VQA and captioning are poor benchmarks for image-to-text ICL. (a) Evaluating state-of-the-art VLLMs on representative examples of popular image-to-text ICL benchmarks – MathVista, VizWiz, and COCO – with standard evaluation protocol. Zero-shot performance is high, and ICL performance depends only weakly on shots, showing that ICL does not learn much. (b) Re-evaluation of VLLMs with LLM-based evaluation further reduces dependence on shots. (c) The impact of ICL on performance goes from small to negligible when moving from traditional to LLM-based evaluation. ICL on these benchmarks primarily learns answer style/format.
@@ 3.1 Main Multimodal Benchmark
Our VL-ICL Bench covers a number of tasks, which includes diverse ICL capabilities such as concept binding, reasoning or fine-grained perception. It covers both image-to-text and text-to-image generation. Our benchmark includes the following eight tasks: Fast Open MiniImageNet, CLEVR Count Induction, Operator Induction, Interleaved Operator Induction, TextOCR, Matching MiniImageNet, Text-to-image MiniImageNet and CoBSAT. We provide illustrations of the tasks in Figure 1, and summarise the diverse capabilities tested by each VL-ICL Bench task in Table 1. This table also summarises the dataset statistics, demonstrating that VL-ICL Bench is compact and accessible. We follow the typical protocol of the ICL community (Dong et al., 2022; Tsimpoukelli et al., 2021; Min et al., 2022) [^2] This is is different than the few-shot meta-learning community (Wang et al., 2020; Hospedales et al., 2021), which samples support/query sets from the same pool. and split each dataset into train and test splits. Few-shot in-context learning is then performed/evaluated by sampling the support/context set from the training split, and the test/query examples from the testing split. Final performance is the average of a number of such ICL episodes.
Table 1: VL-ICL Bench overview. It evaluates diverse capabilities and challenges of ICL with VLLMs. Meanwhile it is compact and easy to be used by researchers, without prohibitive resource requirements.
@@
Total T2I, I2T, Binding, Perception, Long-Context, Interleaving, Induction, Reasoning 14,160 1,520 1.72
@@
Fast Open MiniImageNet  We use the variant of MiniImageNet few-shot object recognition (Vinyals et al., 2016) repurposed for ICL in (Tsimpoukelli et al., 2021). In open-ended classification VLLMs must learn to open-endedly name an object based on a few examples, rather than simply classifying it into a pre-defined set of options. Thus the chance rate is effectively zero, rather than dependent on the number of categories. Fast-binding tasks test the ability of models to associate novel names or symbols to concepts, without relying on prior knowledge. Thus, (Tsimpoukelli et al., 2021) give synthetic names to object categories in the context/support set (e.g. dax or perpo), and the model must learn to associate the names with the illustrated visual concept in order to correctly name the test images. We use the two-way version.
@@
CLEVR Count Induction  In this dataset, models must learn to solve tasks of the type ‘‘How many red objects are there in the scene?’’. However, they must learn this from example rather than being explicitly prompted to do so. Specifically, we input $x$ scene images from CLEVR (Johnson et al., 2017), along with an attribute: value pair that identifies a particular type of object within the scene. There are four types of attributes: size, shape, color, material. The required output $y$ is the count of the objects in the image that meet the provided attribute: value criterion. To solve this task, the model has to provide fine-grained perception to support counting and differentiating object types, learn to ground the requested attribute in images, and importantly induce that the required operator mapping $x$ to $y$ is counting the specified object type [^3] This task could be performed zero-shot with a suitably detailed VQA prompt. However, the goal is to test whether models can learn the task from a few examples by ICL..
@@
Operator Induction  In this image-to-text task, models must solve tasks of the type 2 ? 7 = 9 given training examples like 1 ? 3 = 4. It means that besides parsing an image $x$ to extract the numbers and the operator, models need to induce that the role of the unknown operator is addition, and then conduct simple arithmetic reasoning to apply the induced operator on the parsed test examples. Available mathematical operations are plus, minus and times, and we only consider single digit numbers. We generate our own images for this task.
@@
Interleaved Operator Induction  This task tests the ability of models to reason over multiple images within $x$ to produce a single answer $y$. In this variation of operator induction we give the model two query images as input containing each number in the expression, rather than a single image containing the whole expression, as above. Otherwise it is the same as the basic operator induction task. In one sense, separating the images makes the task easier, as it substantially simplifies the perception task of parsing the expression from a single image. However, it is also harder in that it requires the VLLM to perform induction and reasoning between two different images, when VLLM training (e.g., captioning) typically trains models to produce outputs that depend on a single image at a time. By introducing multiple images, it also introduces a greater number of tokens than basic operator induction, and thus stresses VLLMs’ ability to use a larger context length.
@@
TextOCR  We repurpose the TextOCR dataset (Singh et al., 2021) to create a task where the model should learn to output the text shown in the red rectangle, as inspired by (Sun et al., 2023a). Images $x$ contain an image with a window of text highlighted, and outputs $y$ are the OCR text. This task could be achieved by a suitably detailed zero-shot prompt, but unlike (Sun et al., 2023a) we focus on evaluating whether the task can be induced by way of example through ICL. Thus this task tests both fine-grained perception and induction capabilities.
@@
Matching MiniImageNet  This task is the simplest example of supervised learning of a relation between two images. For relation learning, inputs contain an image pair $x=\{x_{1},x_{2}\}$, and output $y$ indicates whether a specific relation $r$ holds between them. VLLMs are required to learn the relation from examples where it does ($r(x_{1},x_{2})=true$) and does not ($r(x_{1},x_{2})=false$) hold. We re-use MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016) dataset and the relation to learn is whether the two images come from the same class or not (Sung et al., 2018), with disjoint sets of classes considered between training and testing. This task tests induction, the ability to process multiple interleaved images, and the ability to process larger context lengths.
@@
Text-to-Image MiniImageNet  This novel task tests fast concept binding in the text-to-image context. We introduce a further variation of MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016), which inputs synthetic category names $x$ (for fast binding), and outputs images $y$. The model should learn from the context set to associate synthetic names with distributions over images, and thus learn to generate a new image of the corresponding category when prompted with the artificial category name (Figure 1). This task tests image generation and fast binding. LLaVA-Next-7B is used to assess generation correctness.
@@
CoBSAT  We finally also utilize a recent text-to-image CoBSAT (Zeng et al., 2024) benchmark as part of our larger VL-ICL Bench suite (Figure 1). This is a text-to-image task where the model must learn to synthesise images $y$ of a specified text concept $x$ (e.g., object category), but furthermore there is a latent variable common to the context set examples that must be induced and correctly rendered in novel testing images (e.g., common color of objects). This task tests image generation and latent variable induction.
@@
Capability Summary  The VL-ICL Bench suite described above goes far beyond any individual existing ICL benchmark to test diverse capabilities of multimodal ICL including (Table 1): Both text-to-image and image-to-text generation; fast-binding – the ability to rapidly ground new symbols to visual concepts and re-use those symbols in the context of new data; fine-grained perception – as required to count or read text; interleaving – the ability to reason over the content of multiple images when generating a single output; rule induction – inducing non-trivial concepts such as mathematical operators and latent variables from examples; simple reasoning such as arithmetic; and long-context – the ability of a VLLM to usefully exploit a large number of context tokens.
3.2 Text Variation
@@
In order to compare the impact of multimodality, we also include text-version alternatives for our tasks. For datasets such as open-ended MiniImageNet, instead of images we provide image captions and use those for reasoning. For example, in CLEVR we provide enumerations of the objects in the scene, including their attributes. Note that text versions are not practical for all of the tasks, in particular, TextOCR is difficult to translate into a suitable text alternative.

### v2

# VL-ICL Bench: The Devil in the Details of Multimodal In-Context Learning

@@ Abstract
Large language models (LLMs) famously exhibit emergent in-context learning (ICL) – the ability to rapidly adapt to new tasks using few-shot examples provided as a prompt, without updating the model’s weights. Built on top of LLMs, vision large language models (VLLMs) have advanced significantly in areas such as recognition, visual question answering (VQA), reasoning, and grounding. However, investigations into multimodal ICL have predominantly focused on few-shot VQA and image captioning, which we will show neither exploit the strengths of ICL, nor test its limitations. The broader capabilities and limitations of multimodal ICL remain under-explored. In this study, we introduce a comprehensive benchmark for multimodal in-context learning. Our VL-ICL Bench encompasses a broad spectrum of tasks that involve both images and text as inputs and outputs, and different types of challenges, from perception to reasoning and long context length. We evaluate the abilities of state-of-the-art VLLMs on this benchmark suite, revealing their diverse strengths and weaknesses, and showing that even the most advanced models, such as GPT-4, find the tasks challenging. By highlighting a range of new ICL tasks, and the associated strengths and limitations of existing models, we hope that our dataset will inspire future work on enhancing the in-context learning capabilities of VLLMs, as well as inspire new applications that leverage VLLM ICL.
@@ 1 Introduction
With the scaling of model size, large language models (LLMs) famously exhibit the emergent capability of in-context learning (ICL) (Brown et al., 2020; Dong et al., 2022). This refers to the ability to learn from analogy within a single feed-forward pass – thus, enabling the model to learn completely new tasks using a few input-output examples, without requiring any updates to the model parameters. This training-free nature of ICL has led to its rapid and broad application across a wide range of scenarios and applications, as illustrated by benchmarks such as Hendrycks et al. (2021); Zhong et al. (2023); Srivastava et al. (2023); Cobbe et al. (2021).
@@
Vision large language models (VLLMs) are typically built on a base LLM, by augmenting it with a vision encoder and/or decoder connected by some stitching mechanism (Liu et al., 2023a, b; Bai et al., 2023; Zong et al., 2024; Alayrac et al., 2022; Ge et al., 2024). These models have rapidly advanced alongside LLMs, and attracted significant attention for their remarkable multi-modal capabilities in zero-shot recognition, reasoning, grounding, and visual question answering (VQA) among other capabilities. These capabilities have been thoroughly tested by a range of recent benchmark suites (Liu et al., 2023c; Fu et al., 2023; Li et al., 2023c; Yu et al., 2023). Meanwhile, VLLMs are also widely presumed to inherit in-context learning (ICL) capabilities from their base LLM, however, their abilities in this respect are poorly evaluated and poorly understood. Current VLLMs studies mainly report their zero-shot capabilities measured by the benchmarks above, while ICL is usually only evaluated qualitatively, or as a secondary consideration via few-shot visual question answering (VQA) or image captioning (Bai et al., 2023; Awadalla et al., 2023; Sun et al., 2023a; Laurençon et al., 2023), with a notable deficiency in quantitative assessment across a wider spectrum of ICL tasks. This is presumably due to the ready availability of VQA and captioning benchmarking infrastructure. However, we will show that captioning and VQA tasks are not ideal for ICL evaluation: They neither truly exploit the ability of ICL to improve performance from examples; nor do they test the limits of what ICL can do. Thus there is a lack of motivation for future VLLM research to better exploit and expose the underpinning LLM’s ICL ability.
@@
Figure 1: Illustration of the different tasks in VL-ICL Bench. Image-to-text tasks are in the first three rows, while text-to-image tasks are in the bottom two rows. Image-to-text tasks in the third row do reasoning on interleaved image-text inputs.
@@
To enhance the understanding of multimodal ICL and assess the ICL capabilities of state-of-the-art VLLMs, we introduce a novel benchmark suite VL-ICL Bench (Figure 1), tailored for assessing VLLM in-context learning. Our benchmark suite incorporates both text-output and image-output tasks, and is designed to test various facets of VLLMs, including fine-grained perception, reasoning, rule induction, and context-length. We conduct comprehensive evaluations of state-of-the-art VLLMs that are capable of processing interleaved image-text as inputs on our benchmark. Results reveal that although certain models exhibit reasonable performance on specific tasks, none demonstrate uniform excellence across the entire spectrum of tasks, and some models perform near chance level on some tasks. We hope that this systematic study of different opportunities and challenges for multi-modal ICL will support practitioners to know what is currently possible and impossible in terms of training-free learning of new multi-modal tasks, and spur VLLM model developers to study how to expose as much as possible of the LLM’s ICL ability to the multi-modal world.
@@
To summarise our contributions: (1) We demonstrate the limitations inherent in the common practice of quantitatively evaluating VLLM ICL via VQA and captioning. (2) We introduce the first thorough and integrated benchmark suite of ICL tasks covering diverse challenges including perception, reasoning, rule-induction, long context-length and text-to-image/image-to-text. (3) We rigorously evaluate a range of state of the art VLLMs on our benchmark suite, and highlight their diverse strengths and weaknesses, as well the varying maturity of solutions to different ICL challenges.
Figure 2: Performance of top-5 image-to-text models and top-3 text-to-image models on our benchmark. Left: By dataset separately. Right: By capability evaluated, averaging over datasets.
@@ 2.2 Common Practice in ICL Evaluation
The benchmarks that have been most popular in prior attempts at quantitative evaluation of multi-modal ICL are VQA and image captioning. We focus our discussion in this section on image-to-text models (Alayrac et al., 2022; Bai et al., 2023; Li et al., 2023b; Laurençon et al., 2023; Awadalla et al., 2023), as in-context text-to-image models (Ge et al., 2024; Koh et al., 2023) are relatively less common and less mature, so there is no common evaluation practice yet. In the case of captioning, the context set $S$ contains examples of images $x$ and captions $y$; while for VQA the context $S$ contains image-question pairs $x$ and answers $y$.
@@
Figure 3(a) plots the ICL performance of six popular VLLMs on three widely used benchmarks - MathVista VQA (Lu et al., 2024), VizWiz VQA (Gurari et al., 2018), and COCO Captioning (Lin et al., 2014) for varying numbers of training examples (shots). While the performance of the different models varies, the key observation is that most of lines/models show only limited improvement with ICL (shots $>0$) compared to the zero-shot case (shots $=0$). This is because, while the context set $S$ illustrates the notion of asking and answering a question or captioning images, the baseline VLLM $\theta$ is already quite good at VQA and captioning. The limiting factors in VLLM captioning and VQA are things like detailed perception, common sense knowledge, etc. – all of which are fundamental challenges to the VLLM, and not things that can reasonably be taught by a few-shot support set.
@@
Given the discussion above, it is unclear why performance should improve with shots at all? We conjectured that this is largely due to the VLLM learning about each dataset’s preferred answer style, rather than learning to better solve multi-modal inference tasks per-se. For example, in captioning zero-shot VLLMs tend to produce more verbose captions than COCO ground-truth, and they learn to be more concise through ICL. Meanwhile, for VQA, there is a standard practice of evaluating based on string match between the ground-truth answer and the model-provided answer. For example, VizWiz has unanswerable questions, which some VLLMs answer with "I don’t know" which would not be string matched against a ground truth "Unanswerable". Some models thus learn about answer-formatting (e.g., preferred terminology; avoid using any preface or postface that may throw of a string match) from the context set. This is indeed a kind of ICL, but perhaps not what one expects to be learning in VQA. To validate this conjecture, we repeat the previous evaluation, but using soft matching to eliminate the impact of answer format learning. For VQA, we use a pretrained LLM to determine whether the prediction has semantically equivalent to the ground-truth while for captioning, we use GPT-3.5 to score the quality of the generated caption on a scale of score 1-10 (details in appendix). Fig. 3(b) shows that the curves have almost fully flattened out, with zero-shot performance having improved. Fig. 3(c) quantifies this difference by showing the average rate of improvement with shots for exact match and LLM match. The change to LLM validation almost completely eliminates any benefit of ICL over zero-shot.
@@
In contrast to the above, popular LLM ICL benchmarks in the lanugage domain do usually exhibit non-trivial ICL learning (Brown et al., 2020; Dong et al., 2022). Figure 4 shows three state of the art VLLMs along with their corresponding base LLMs, evaluated on three popular NLP tasks (AGNews (Zhang et al., 2015), MIT Movies (Ushio and Camacho-Collados, 2021) and TREC (Voorhees and Tice, 2000)). We can see that in contrast to the VQA/Captioning benchmarks, models’ zero-shot performance is often substantially improved by few-shot ICL. This result confirms that the LLM components in VLLMs do inherit the ICL ability of their base LLM. However, it raises the question of how we can meaningfully exploit and measure the ICL ability of VLLMs in the multi-modal context. In the next section, we introduce our benchmark VL-ICL Bench, which does exactly this.
(a) Exact match evaluation.
@@
Figure 3: VQA and Captioning are poor benchmarks for image-to-text ICL. (a) Evaluating state-of-the-art VLLMs on representative examples of popular image-to-text ICL benchmarks - MathVista, VizWiz, and COCO - with standard evaluation protocol. Zero-shot performance is high, and ICL performance depends only weakly on shots, showing that ICL does not learn much. (b) Re-evaluation of VLLMs with LLM-based evaluation further reduces dependence on shots. (c) The impact of ICL on performance goes from small to negligible when moving from traditional to LLM-based evaluation. ICL on these benchmarks primarily learns answer style/format.
@@ 3.1 Main Multimodal Benchmark
Our VL-ICL Bench covers a number of tasks, which includes diverse ICL capabilities spanning concept binding, reasoning or fine-grained perception. It covers both image-to-text and text-to-image generation. Our benchmark includes ten tasks detailed below, illustrated in Figure 1. Table 1 summarises the diverse capabilities tested by each VL-ICL Bench task, and demonstrates its compactness. We follow the typical protocol of the ICL community (Dong et al., 2022; Tsimpoukelli et al., 2021; Min et al., 2022) [^2] This is is different than the few-shot meta-learning community (Wang et al., 2020; Hospedales et al., 2021), which samples support/query sets from the same pool. and split each dataset into train and test splits. Few-shot ICL is then performed/evaluated by sampling the support/context set from the training split, and the test/query examples from testing split. The final performance is the average of a number of such ICL episodes.
Table 1: VL-ICL Bench overview. It evaluates diverse capabilities and challenges of ICL with VLLMs, while being compact and easy to be use without prohibitive resource requirements.
@@
Fast Counting T2I, Fast Binding 0,800 040 0.03
Fast Attribute Matching T2I, Fast Binding 0,300 200 0.07
@@
Total T2I, I2T, Binding, Perception, Long-Context, Interleaving, Induction, Reasoning 15,260 1,760 1.82
@@
Fast Open MiniImageNet  We use the variant of MiniImageNet few-shot object recognition (Vinyals et al., 2016) repurposed for ICL in Tsimpoukelli et al. (2021). In open-ended classification, VLLMs must name objects based on a few examples, rather than classifying them into predefined categories, making the chance rate effectively zero. Fast-binding tasks test models’ ability to associate novel names or symbols with concepts, without prior knowledge. Tsimpoukelli et al. (2021) assign synthetic names (e.g., dax or perpo) to object categories, and the model must learn these associations to name test images. We use the two-way version of this task.
@@
CLEVR Count Induction  In this dataset, models must learn to solve tasks like "How many red objects are there in the scene?" from examples rather than explicit prompts. We input CLEVR scene images (Johnson et al., 2017) along with an attribute: value pair identifying a specific object type based on four attributes: size, shape, color, or material. The output $y$ is the count of objects matching the attribute. To succeed, the model must ground the attribute in the image, differentiate object types, and induce that the required operation is counting. [^3] This task could be performed zero-shot with a suitably detailed VQA prompt. But the goal is to test whether models can learn the task from a few examples by ICL..
@@
Operator Induction  In this image-to-text task, models must solve tasks of the type 2 ? 7 = 9 given training examples like 1 ? 3 = 4. I.e., besides parsing an image $x$ to extract the numbers and operator, models need to induce that the role of the unknown operator is addition, and then conduct arithmetic reasoning on parsed test examples. Available mathematical operations are plus, minus and times, and we only consider single digit numbers. We generate our own images for this task.
@@
Interleaved Operator Induction  This task tests the ability of models to reason over multiple images within $x$ to produce a single answer $y$. In this variation of operator induction we give the model two query images as input containing each number in the expression, rather than a single image containing the whole expression, as above. While separating the images simplifies perception difficulty of parsing expressions, it also increases difficulty by requiring reasoning between two images, whereas VLLMs are typically trained on single-image tasks. Additionally, multiple images increase the token count, challenging the VLLMs’ ability to handle larger context lengths.
@@
TextOCR  We repurpose the TextOCR dataset (Singh et al., 2021) to create a task where the model should learn to output the text shown in the red rectangle, as inspired by (Sun et al., 2023a). Images $x$ contain an image with a window of text highlighted, and outputs $y$ are the OCR text. This task could be achieved by a suitably detailed zero-shot prompt, but unlike Sun et al. (2023a), we focus on evaluating whether the task can be induced by way of example through ICL. Thus this task tests both fine-grained perception and induction capabilities.
@@
Fast Matching MiniImageNet  This task is the simplest example of supervised learning of a relation between two images. For relation learning, inputs contain an image pair $x=\{x_{1},x_{2}\}$, and output $y$ indicates whether a specific relation $r$ holds between them. VLLMs are required to learn the relation from examples where it does ($r(x_{1},x_{2})=true$) and does not ($r(x_{1},x_{2})=false$) hold, using artificial keywords to name the relation. We re-use MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016) and the relation to learn is whether the two images come from the same class or not (Sung et al., 2018). This task tests induction, long-context, and multiple interleaved images.
@@
Fast Text-to-Image MiniImageNet  We introduce a further variation of MiniImageNet (Tsimpoukelli et al., 2021; Vinyals et al., 2016), which inputs synthetic category names $x$ (for fast binding), and outputs images $y$. The model should learn from the context set to associate synthetic names with distributions over images, and thus learn to generate a new image of the corresponding category when prompted with the artificial category name. This task tests image generation and fast binding.
@@
CoBSAT  We also utilize a recent text-to-image CoBSAT (Zeng et al., 2024) benchmark as part of our larger VL-ICL Bench suite (Figure 1). This is a text-to-image task where the model must learn to synthesise images $y$ of a specified text concept $x$ (e.g., object category), but furthermore there is a latent variable common to the context set examples that must be induced and correctly rendered in novel testing images (e.g., common color of objects). This task tests image generation and latent variable induction. We also use selected CoBSAT images for Fast Counting and Fast Attribute tasks.
Fast Counting  In the fast counting task we associate artificial names with counts of objects in the image. The task is to generate an image that shows a given object in quantity associated with the keyword (e.g. perpo dogs where perpo means two). We use counts between one and four to make the task solvable, and ask the model to distinguish between two counts in a task.
Fast Attribute Matching  Similarly to the other tasks, we ask the model to learn to associate artificial names to a concept, in this task to attributes (e.g. color red). We then ask the model to generate an image of an object with the given attribute, when given the name of the object and the keyword associated with the attribute. We show the model two attributes in the support set.
@@
Capability Summary  The VL-ICL Bench suite described above goes far beyond any individual existing ICL benchmark to test diverse capabilities of multi-modal ICL including (Table 1): Both text-to-image and image-to-text generation; fast-binding – the ability to rapidly ground new symbols to visual concepts and re-use those symbols in the context of new data; fine-grained perception – as required to count or read text; interleaving – the ability to reason over the content of multiple images when generating a single output; rule induction – inducing non-trivial concepts such as mathematical operators and latent variables from examples; simple reasoning such as arithmetic; and long-context – the ability of a VLLM to usefully exploit a large number of context tokens.
@@
Text Variation  In order to compare the impact of multimodality we also include text-version alternatives for our tasks. For datasets such as open-ended MiniImageNet, instead of images we provide image captions and use those for reasoning. For example, in CLEVR we provide enumeration of the objects in the scene, including their attributes. Note that text versions are not practical for all of the tasks, in particular TextOCR is difficult to translate into a suitable text alternative.